OpenStudy (anonymous):

Which set of side lengths defines a right triangle? A. 4, 8, 9 B. 5, 9, 10 C. 6, 10, 13 D. 7, 24, 25

OpenStudy (anonymous):

b i think

OpenStudy (anonymous):

but u need angles to define that because A right triangle (American English) or right-angled triangle is a triangle in which one angle is a right angle (that is, a 90-degree angle).

OpenStudy (unknownunknown):

Do Sarah, all you need to do here is apply the pythagorean theorem. But I'm going to introduce you to something you have not learned yet. We can calculate any side in a triangle with the formula, c^2 = a^2 + b^2 - cos(theta)*2*a*b Now, if we have a 90 degree triangle, theta becomes 90 degrees, and cos(theta) = 0. Therefore, this simplifies to c^2 = a^2 + b^2, which is the pythagorean theorem you're familiar with (but take note of the more expanded formula you can apply to a triangle of any angle, even those that aren't right triangles!). So, looking at D. The square root of 7^2 + 24^2 is what?

OpenStudy (unknownunknown):

\[\sqrt{7^{2} + 24^{2}} = ?\]

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (unknownunknown):

No problem =)